# TLDR: Unsupervised Goal-Conditioned RL via Temporal Distance-Aware Representations

### v1

@@ Abstract
Unsupervised goal-conditioned reinforcement learning (GCRL) is a promising paradigm for developing diverse robotic skills without external supervision. However, existing unsupervised GCRL methods often struggle to cover a wide range of states in complex environments due to their limited exploration and sparse or noisy rewards for GCRL. To overcome these challenges, we propose a novel unsupervised GCRL method that leverages TemporaL Distance-aware Representations (TLDR). TLDR selects faraway goals to initiate exploration and computes intrinsic exploration rewards and goal-reaching rewards, based on temporal distance. Specifically, our exploration policy seeks states with large temporal distances (i.e. covering a large state space), while the goal-conditioned policy learns to minimize the temporal distance to the goal (i.e. reaching the goal). Our experimental results in six simulated robotic locomotion environments demonstrate that our method significantly outperforms previous unsupervised GCRL methods in achieving a wide variety of states.
@@ 1 Introduction
Figure 1: Trajectories (red) of an ant robot in a complex maze trained by TLDR (ours), METRA [1], and PEG [2]. While prior methods yield limited exploration, TLDR explores the entire maze.
@@
The major challenges in unsupervised GCRL are twofold: (1) exploring diverse states to ensure the agent can learn to achieve a wide variety of goals, and (2) effectively learning a goal-reaching policy. Previous methods focus on exploring novel states [11] or states with high uncertainty in next state prediction [10, 2]. However, these methods aim to discover unseen states or state transitions, which may not be meaningful. Additionally, training a goal-reaching policy to maximize sparse goal-reaching rewards [8] or minimize heuristically-defined distances to a goal [10, 12] is often insufficient for long-horizon goal-reaching behaviors in complex environments.
@@
TLDR demonstrates superior state coverage compared to prior unsupervised GCRL and skill discovery methods in complex AntMaze environments (see Figure 1). Our ablation studies confirm that our temporal distance-aware approach enhances both goal-directed exploration and goal-conditioned policy learning. Furthermore, our method outperforms prior work across diverse locomotion environments, underscoring its general applicability.
@@ 2 Related Work
Unsupervised goal-conditioned reinforcement learning (GCRL) aims to learn a goal-conditioned policy that can reach diverse goal states without external supervision [15, 16, 10, 2]. The major challenges of unsupervised GCRL can be summarized in two aspects: (1) optimizing goal-conditioned policies and (2) collecting trajectories with novel goals that effectively enlarge its state coverage.
@@
Recent techniques such as hindsight experience reply (HER) [8] and model-based policy optimization [10, 12] have improved the efficiency of GCRL. However, learning complex, long-horizon goal-reaching behaviors remains difficult due to sparse (e.g. whether it reaches the goal) [8] or heuristic rewards (e.g. cosine similarity between the state and goal) [10, 12]. Instead, temporal distance, defined as the number of environment steps between states estimated from data, can provide more dense and grounded rewards [17, 10, 18]. Nonetheless, this often leads to sub-optimal goal-reaching behaviors since it does not reflect the “shortest temporal distance” between states. In this paper, we propose to use the estimated shortest temporal distance as reward signals for GCRL, inspired by QRL [14] and HILP [13]. We apply the learned representations to compute goal-reaching rewards rather than directly learning the value function in QRL or using it for skill-learning rewards in HILP.
@@
Exploration in unsupervised GCRL relies heavily on selecting exploratory goals that lead to novel states and expand state coverage. Various strategies for exploratory goal selection have been introduced, including selecting less visited states [19], states with low-density in state distributions [20, 11], and states with high uncertainty in dynamics [10, 2]. Instead of sampling uncertain or less visited states as goals, we select goals that are temporally distant from the visited state distribution, encouraging coverage of broader state spaces which require more environment steps to reach.
@@
Unsupervised skill discovery [21, 22, 23, 24, 25, 26, 27, 1] is another approach to learning diverse behaviors without supervision, yet often lacks robust exploration capabilities [27], requiring manual feature engineering or limiting to low-dimensional state spaces. METRA [1] addresses these limitations by computing skill-learning rewards with temporal distance-aware representations, though it exhibits limited coverage in complex environments, as depicted in Figure 1.
Temporal distance-aware representations have been extensively used in imitation learning [28], representation learning [29, 30], unsupervised skill discovery [1], offline skill learning [13], and GCRL [14]. Methods like QRL [14], HILP [13], and METRA [1] are closely related to our work as they learn temporal distance-preserving representations or goal-reaching value functions. HILP and METRA use temporal distance-aware representations for skill representations and skill rewards. On the other hand, QRL learns a quasimetric model as a (negated) goal-reaching value function. In contrast, our method uses temporal distance-aware representations across the entire unsupervised GCRL pipeline.
@@
Figure 2: Overview of TLDR algorithm. TLDR leverages temporal distance-aware representations for unsupervised GCRL (Go-Explore [19] in this paper). (a) We start by learning a state encoder $\phi$ that maps states to temporal distance-aware representations. With the temporal distance-aware representations, TLDR (b) selects the temporally farthest state from the visited states as an exploratory goal, (c) reaches the chosen goal using a goal-conditioned policy, which learns to minimize temporal distance to the goal, and (d) collects exploratory trajectories using an exploration policy that visits states with large temporal distance from the visited states.
@@ 3 Approach
In this paper, we introduce TemporaL Distance-aware Representations (TLDR), an unsupervised goal-conditioned reinforcement learning (GCRL) method, integrating temporal distance within unsupervised GCRL. As illustrated in Figure 2, TLDR integrates temporal distance-aware representations (Section 3.2) into every facet of the Go-Explore [19] strategy (Section 3.3), which chooses a goal from experience (Section 3.4), reaches the selected goal via the goal-conditioned policy, and executes the exploration policy to gather diverse experiences. We then refine both the exploration policy (Section 3.5) and goal-conditioned policy (Section 3.6) based on the collected data and rewards computed using the temporal distance-aware representations. We describe the full algorithm in Algorithm 1. Please refer to Appendix A for further implementation details.
@@ 3.1 Problem Formulation
We formulate the unsupervised GCRL problem with a goal-conditioned Markov decision process, defined as the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},p,\mathcal{G})$. $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, respectively. $p:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ denotes the transition dynamics, where $\Delta(\mathcal{X})$ denotes the set of probability distributions over $\mathcal{X}$. The goal of the agent is to learn an optimal goal-conditioned policy $\pi^{G}:\mathcal{S}\times\mathcal{G}\rightarrow\mathcal{A}$, where $\pi^{G}(\mathbf{a}\mid\mathbf{s},\mathbf{g})$ outputs an action $\mathbf{a}\in\mathcal{A}$ that can navigate to the goal $\mathbf{g}\in\mathcal{G}$ as fast as possible from the current state $\mathbf{s}$. In this paper, we set $\mathcal{G}=\mathcal{S}$, allowing any state as a potential goal for the agent.
@@ 3.2 Learning Temporal Distance-Aware Representations
Temporal distance, defined as the minimum number of environment steps between states, can provide more dense and grounded rewards for goal-conditioned policy learning as well as exploration. For GCRL, instead of relying on sparse and binary goal-reaching rewards, the change in temporal distance before and after taking an action can be an informative learning signal. Moreover, exploration in unsupervised GCRL can be incentivized by discovering temporally faraway states. Therefore, in this paper, we propose to use temporal distance for unsupervised GCRL.
We first estimate the temporal distance by learning temporal distance-aware representations, inspired by Park et al. [13], Wang et al. [14]. We learn the representation $\phi:\mathcal{S}\rightarrow\mathcal{Z}$, that encodes the temporal distance between two states into the latent space $\mathcal{Z}$, where $\lVert\phi(\mathbf{s}_{1})-\phi(\mathbf{s}_{2})\rVert$ represents the shortest temporal distance between $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$. This representation is then used across the entire unsupervised GCRL algorithm: exploratory goal selection, intrinsic reward for exploration, and reward for goal-conditioned policy.
To train temporal distance-aware representations, we adopt a constrained optimization approach similar to the objective of QRL [14]:
$\ \max_{\phi}\mathbb{E}_{\mathbf{s}\sim p_{\mathbf{s}},\mathbf{g}\sim p_{\mathbf{g}}}\left[\lVert f(\phi(\mathbf{s})-\phi(\mathbf{g}))\rVert\right]\quad\text{ s.t. }\quad\mathbb{E}_{(\mathbf{s},\mathbf{a},\mathbf{s}^{\prime})\sim p_{\text{transition}}}\left[\lVert\phi(\mathbf{s})-\phi(\mathbf{s}^{\prime})\rVert\right]\leq 1,$ (1)
where $f$ represents an affine-transformed softplus function that assigns lower weights to larger distances $\lVert\phi(\mathbf{s})-\phi(\mathbf{g})\rVert$. We optimize this constrained objective using dual gradient descent with a Lagrange multiplier $\lambda$, and we randomly sample $\mathbf{s}$ and $\mathbf{g}$ from a minibatch during training.
@@ 3.3 Unsupervised GCRL with Temporal Distance-Aware Representations
With temporal distance-aware representations, we can integrate the concept of temporal distance into unsupervised GCRL. Our approach is built upon the Go-Explore procedure [19], a widely-used unsupervised GCRL algorithm comprising two phases: (1) the “Go-phase,” where the goal-conditioned policy $\pi^{G}$ navigates toward a goal $\mathbf{g}$, and (2) the “Explore-phase,” where the exploration policy $\pi^{E}$ gathers new state trajectories to refine the goal-conditioned policy.
While Go-Explore relies on task-specific information for goal selection and exploration policy training, our method uses task-agnostic temporal distance metrics induced by temporal distance-aware representations. The subsequent sections detail how our method leverages the representation for selecting goals in the Go-phase (Section 3.4), enhancing the exploration policy (Section 3.5), and facilitating the GCRL policy training (Section 3.6).
@@
1:  Initialize goal-conditioned policy $\pi^{G}_{\theta}$, exploration policy $\pi^{E}_{\theta}$, and temporal distance-aware representation $\phi$
2:  $\mathcal{D}\leftarrow\emptyset$
3:  while not converged do
4:     $\mathbf{s}_{0}\sim p(\mathbf{s}_{0})$
5:     Sample a minibatch $\mathcal{B}\sim\mathcal{D}$
6:     $\mathbf{g}\leftarrow\arg\max_{\mathbf{s}\in\mathcal{B}}(r_{\text{TLDR}}(\mathbf{s}))$triangleright Select state with the highest TLDR reward (Eq. 2)
7:     for $t=0,\ldots,T-1$ do
8:        if $t<T_{G}$ then
9:           $\mathbf{a}_{t}\sim\pi^{G}_{\theta}(\cdot\mid\mathbf{s}_{t},\mathbf{g})$triangleright Follow goal-conditioned policy $\pi^{G}_{\theta}$ for $T_{G}$ steps
10:        else
11:           $\mathbf{a}_{t}\sim\pi^{E}_{\theta}(\cdot\mid\mathbf{s}_{t})$triangleright Explore using exploration policy $\pi^{E}_{\theta}$
12:        $\mathbf{s}_{t+1}\sim p(\cdot\mid\mathbf{s}_{t},\mathbf{a}_{t})$
13:        $\mathcal{D}\leftarrow\mathcal{D}\cup\{\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{s}_{t+1}\}$
@@ 3.4 Exploratory Goal Selection
For unsupervised GCRL, selecting low-density (less visited) states as exploratory goals can enhance goal-directed exploration [15, 16]. However, the concept of “density” of a state does not necessarily indicate how rare or hard to reach the state. For example, while a robotic arm might actively seek out unseen (low-density) joint positions, interacting with objects could offer more significant learning opportunities [27]. Thus, we propose selecting goals that are temporally distant from states that are already visited (i.e. in the replay buffer) to explore not only diverse but also hard-to-reach states.
To sample a faraway goal at the start of each episode, we employ the non-parametric particle-based entropy estimator [25] on top of our temporal distance-aware representations. We choose $N$ goals with the top-$N$ highest entropy, which we refer to as TLDR reward, and collect $N$ corresponding trajectories using the goal-reaching policy. The TLDR reward for each state is computed as follows:
@@
where $N_{k}(\cdot)$ denotes the $k$-nearest neighbors around $\phi(\mathbf{s})$ within a single minibatch.

### v2

@@ Abstract
Unsupervised goal-conditioned reinforcement learning (GCRL) is a promising paradigm for developing diverse robotic skills without external supervision. However, existing unsupervised GCRL methods often struggle to cover a wide range of states in complex environments due to their limited exploration and sparse or noisy rewards for GCRL. To overcome these challenges, we propose a novel unsupervised GCRL method that leverages TemporaL Distance-aware Representations (TLDR). Based on temporal distance, TLDR selects faraway goals to initiate exploration and computes intrinsic exploration rewards and goal-reaching rewards. Specifically, our exploration policy seeks states with large temporal distances (i.e. covering a large state space), while the goal-conditioned policy learns to minimize the temporal distance to the goal (i.e. reaching the goal). Our results in six simulated locomotion environments demonstrate that TLDR significantly outperforms prior unsupervised GCRL methods in achieving a wide range of states.
@@ 1 Introduction
Figure 1: Trajectories (red) of an ant robot in a complex maze trained by TLDR, METRA [1], and PEG [2]. While prior methods yield limited exploration, TLDR explores the entire maze.
@@
The major challenges in unsupervised GCRL are twofold: (1) exploring diverse states that the agent can learn to achieve, and (2) effectively learning a goal-reaching policy. Prior unsupervised GCRL methods focus on exploring novel states [11] or states with high uncertainty in next state prediction [10, 2]. However, discovering unseen states or state transitions may not lead to meaningful states. Additionally, training a goal-reaching policy to maximize sparse [8] or heuristic [10, 12] goal-reaching rewards is often insufficient for long-horizon goal-reaching behaviors in complex environments.
@@
TLDR demonstrates superior state coverage compared to prior unsupervised GCRL and skill discovery methods in complex AntMaze environments, as shown in Figure 1. Our ablation studies confirm that our temporal distance-aware approach enhances both goal-directed exploration and goal-conditioned policy learning. Furthermore, our method outperforms prior work across diverse locomotion environments, underscoring its general applicability.
@@ 2 Related Work
Unsupervised goal-conditioned reinforcement learning (GCRL) aims to learn a goal-conditioned policy that can reach diverse goal states without external supervision [15, 16, 10, 2]. The major challenges of unsupervised GCRL can be summarized in two aspects: (1) optimizing a goal-conditioned policy and (2) collecting trajectories with novel goals that effectively enlarge its state coverage.
To improve the efficiency of goal-conditioned policy learning, hindsight experience reply (HER) [8] and model-based policy optimization [10, 12] have been widely used. However, learning complex, long-horizon goal-reaching behaviors remains difficult due to sparse (e.g. whether it reaches the goal [8]) or heuristic rewards (e.g. cosine similarity between the state and goal [10, 12]).
@@
Instead, temporal distance, defined as the number of environment steps between states estimated from data, can provide more dense and grounded rewards [17, 10, 18, 19]. LEXA [10] and PEG [2] use the expected temporal distances regarding the current policy as goal-reaching rewards [19]. However, this does not reflect the “shortest temporal distance” between states, often leading to sub-optimal goal-reaching behaviors. In this paper, we propose to use the estimated shortest temporal distance as reward signals for GCRL, inspired by QRL [14] and HILP [13]. We apply the learned representations to compute goal-reaching rewards rather than directly learning the value function in QRL or using it for skill-learning rewards in HILP.
@@
Exploration in unsupervised GCRL relies heavily on selecting exploratory goals that lead an agent to novel states and expand state coverage. Exploratory goals can be simply sampled from a replay buffer as in LEXA [10], or can be selected from less visited states [20], states with low-density in state distributions [21, 11], and states with high uncertainty in dynamics [2]. Instead of sampling uncertain or less visited states as goals, we select states temporally distant from the visited state distribution as goals, encouraging the discovery of temporally farther away states.
In addition to exploratory goal selection, an explicit exploration policy [22, 20] can further encourage exploration by maximizing intrinsic rewards, such as uncertainty in dynamics used in LEXA and PEG. For better exploration, our approach opts for maximizing temporal distance from the visited states, continuously seeking novel and faraway states.
@@
Unsupervised skill discovery [23, 24, 25, 26, 27, 28, 29, 1] is another approach to learning diverse behaviors without supervision, yet often lacks robust exploration capabilities [29], requiring manual feature engineering or limiting to low-dimensional state spaces. METRA [1] addresses these limitations by computing skill-learning rewards with temporal distance-aware representations. While achieving remarkable exploration and zero-shot goal-reaching capabilities, METRA exhibits limited coverage in complex environments, as depicted in Figure 1. We find that METRA tends to focus on reaching the known farthest states rather than exploring less visited states, whereas our exploration strategy explicitly encourages reaching unseen farther states.
@@
Figure 2: Overview of TLDR algorithm. TLDR leverages temporal distance-aware representations for unsupervised GCRL. (a) We start by learning a state encoder $\phi(\mathbf{s})$ that maps states to temporal distance-aware representations. With the temporal distance-aware representations, TLDR (b) selects the temporally farthest state from the visited states as an exploratory goal, (c) reaches the chosen goal using a goal-conditioned policy, which learns to minimize temporal distance to the goal, and (d) collects exploratory trajectories using an exploration policy that visits states with large temporal distance from the visited states.
@@ 3 Approach
In this paper, we introduce TemporaL Distance-aware Representations (TLDR), an unsupervised goal-conditioned reinforcement learning (GCRL) method, integrating temporal distance-aware representations (Section 3.2) into every facet of the Go-Explore strategy [20] (Section 3.3), as illustrated in Figure 2. TLDR first chooses a goal from experience (Section 3.4), reaches the selected goal via the goal-conditioned policy, and executes the exploration policy to gather diverse experiences. Both the exploration policy (Section 3.5) and goal-conditioned policy (Section 3.6) are then trained on the collected data and rewards computed using the temporal distance-aware representations. We describe the full algorithm in Algorithm 1 and implementation details in Appendix A.
@@ 3.1 Problem Formulation
We formulate the unsupervised GCRL problem with a goal-conditioned Markov decision process, defined as the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},p,\mathcal{G})$. $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, respectively. $p:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ denotes the transition dynamics, where $\Delta(\mathcal{X})$ denotes the set of probability distributions over $\mathcal{X}$. The goal of the agent is to learn an optimal goal-conditioned policy $\pi^{G}:\mathcal{S}\times\mathcal{G}\rightarrow\mathcal{A}$, where $\pi^{G}(\mathbf{a}\mid\mathbf{s},\mathbf{g})$ outputs an action $\mathbf{a}\in\mathcal{A}$ that can navigate to the goal $\mathbf{g}\in\mathcal{G}$e state $\mathbf{s}$ within minimum steps. In this paper, we set $\mathcal{G}=\mathcal{S}$, allowing any state as a potential goal for the agent.
@@ 3.2 Learning Temporal Distance-Aware Representations
Temporal distance, defined as the minimum number of environment steps between states, can provide more dense and grounded rewards for goal-conditioned policy learning as well as exploration. For GCRL, instead of relying on sparse and binary goal-reaching rewards, the change in temporal distance before and after taking an action can be an informative learning signal. Moreover, exploration in unsupervised GCRL can be incentivized by discovering temporally faraway states.
Therefore, in this paper, we propose to use temporal distance for unsupervised GCRL. We first estimate the temporal distance by learning temporal distance-aware representations, inspired by Park et al. [13], Wang et al. [14]. The learned representation $\phi:\mathcal{S}\rightarrow\mathcal{Z}$ encodes the temporal distance between two states into the latent space $\mathcal{Z}$, where $\lVert\phi(\mathbf{s}_{1})-\phi(\mathbf{s}_{2})\rVert$ represents the temporal distance between $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$. This representation is then used across the entire unsupervised GCRL algorithm: exploratory goal selection, intrinsic reward for exploration, and reward for a goal-conditioned policy.
To train temporal distance-aware representations, we adopt QRL’s constrained optimization [14]:
$\ \max_{\phi}\mathbb{E}_{\mathbf{s}\sim p_{\mathbf{s}},\mathbf{g}\sim p_{\mathbf{g}}}\left[f(\lVert\phi(\mathbf{s})-\phi(\mathbf{g})\rVert)\right]\quad\text{ s.t. }\quad\mathbb{E}_{(\mathbf{s},\mathbf{a},\mathbf{s}^{\prime})\sim p_{\text{transition}}}\left[\lVert\phi(\mathbf{s})-\phi(\mathbf{s}^{\prime})\rVert\right]\leq 1,$ (1)
where $f$ is an affine-transformed softplus function that assigns lower weights to larger distances $\lVert\phi(\mathbf{s})-\phi(\mathbf{g})\rVert$. We optimize this constrained objective using dual gradient descent with a Lagrange multiplier $\lambda$, and we randomly sample $\mathbf{s}$ and $\mathbf{g}$ from a minibatch during training.
@@ 3.3 Unsupervised GCRL with Temporal Distance-Aware Representations
With temporal distance-aware representations, we can integrate the concept of temporal distance into unsupervised GCRL. Our approach is built upon the Go-Explore procedure [20], a widely-used unsupervised GCRL algorithm comprising two phases: (1) the “Go-phase,” where the goal-conditioned policy $\pi^{G}(\mathbf{a}\mid\mathbf{s},\mathbf{g})$ navigates toward a goal $\mathbf{g}$, and (2) the “Explore-phase,” where the exploration policy $\pi^{E}(\mathbf{a}\mid\mathbf{s})$ gathers new state trajectories to refine the goal-conditioned policy.
While Go-Explore relies on task-specific information for goal selection and executes random actions for exploration, our method uses task-agnostic temporal distance metrics induced by temporal distance-aware representations. The subsequent sections detail how our method leverages the temporal distance-aware representations for selecting goals in the Go-phase (Section 3.4), enhancing the exploration policy (Section 3.5), and facilitating the GCRL policy training (Section 3.6).
@@
1:  Initialize goal-conditioned policy $\pi^{G}_{\theta}$, exploration policy $\pi^{E}_{\theta}$, temporal distance-aware representation $\phi$, and replay buffer $\mathcal{D}$
2:  while not converged do
3:     $\mathbf{s}_{0}\sim p(\mathbf{s}_{0})$
4:     Sample a minibatch $\mathcal{B}\sim\mathcal{D}$
5:     $\mathbf{g}\leftarrow\arg\max_{\mathbf{s}\in\mathcal{B}}(r_{\text{TLDR}}(\mathbf{s}))$triangleright Select state with the highest TLDR reward (Eq. 2)
6:     for $t=0,\ldots,T-1$ do
7:        if $t<T_{G}$ then
8:           $\mathbf{a}_{t}\sim\pi^{G}_{\theta}(\cdot\mid\mathbf{s}_{t},\mathbf{g})$triangleright Follow goal-conditioned policy $\pi^{G}_{\theta}$ for $T_{G}$ steps
9:        else
10:           $\mathbf{a}_{t}\sim\pi^{E}_{\theta}(\cdot\mid\mathbf{s}_{t})$triangleright Explore using exploration policy $\pi^{E}_{\theta}$
11:        $\mathbf{s}_{t+1}\sim p(\cdot\mid\mathbf{s}_{t},\mathbf{a}_{t})$
12:        $\mathcal{D}\leftarrow\mathcal{D}\cup\{\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{s}_{t+1}\}$
13:     Train representations $\phi$ to minimize $\mathcal{L}_{\phi}$ in Eq. 1
@@ 3.4 Exploratory Goal Selection
For unsupervised GCRL, selecting low-density (less visited) states as exploratory goals can enhance goal-directed exploration [15, 16]. However, the concept of the “density” of a state does not necessarily indicate how rare or hard it is to reach the state. For example, while a robotic arm might actively seek out unseen (low-density) joint positions, interacting with objects could offer more significant learning opportunities [29]. Thus, we propose selecting goals that are temporally distant from states that are already visited (i.e. in the replay buffer) to explore not only diverse but also hard-to-reach states.
To sample a faraway goal at the start of each episode, we employ the non-parametric particle-based entropy estimator [27] on top of our temporal distance-aware representations. Among states in a minibatch, we choose $N$ goals with h entropy and collect $N$ corresponding trajectories using the goal-reaching policy. The entropy can be estimated as follows, which we refer to as TLDR reward:
@@
where $N_{k}(\cdot)$ denotes the $k$-nearest neighbors around $\phi(\mathbf{s})$ within a minibatch.